OpenStudy (anonymous):

Can someone help me. Solve the equation: x/6 - 2/3 = x/9

OpenStudy (anonymous):

i would first multiplay both sides by 9 to get rid of the fraction, then multiply by 6 to get rid of the other fractions and make it easier to slove

OpenStudy (yttrium):

Better get rid of the fractions first. If you know what @BigTeach2345 is trying to say. Good. If not, just tell us. :)

OpenStudy (anonymous):

I still don't understand

OpenStudy (skullpatrol):

You can use the idea of balance to solve equations...

OpenStudy (anonymous):

(x/6 -2/3 = x/9) X 9 = (9x/6 - 6 = x) = 9x - 36 = 6x Now you should be able to slove this equation

OpenStudy (skullpatrol):

...whatever you do to one side of the equation you must do it to the other to keep the balance.

OpenStudy (skullpatrol):

$$\huge (\frac{x}{6} -\frac{2}{3})*9 = (\frac{x}{9}) * 9$$ $$\huge \frac{9x}{6} - 6 = x$$ $$\huge9x - 36 = 6x $$

OpenStudy (skullpatrol):

@tryder have you been taught these transformations?